Required information

Use the following information for the Quick Study below.

[The following information applies to the questions displayed below.]

The Carlberg Company has two manufacturing departments, assembly and painting. The assembly department started 10,300 units during November. The following production activity unit and cost information refers to the assembly department’s November production activities.

Assembly Department Units Percent of Direct
Materials Added
Percent of
Conversion
Beginning work in process 3,000 60 % 40 %
Units transferred out 10,000 100 % 100 %
Ending work in process 3,300 80 % 30 %

Beginning work in process
inventory—Assembly dept
$ 2,696 (includes $896 for direct materials
and $1,800 for conversion)
Costs added during the month:
Direct materials $ 13,008
Conversion $ 14,685

rev: 04_18_2018_QC_CS-124631

QS 3-13 Weighted average: Journal entry to transfer costs LO P4

Required:
Prepare the November 30 journal entry to record the transfer of units (and costs) from the assembly department to the painting department. Use the weighted-average method.

  • Record the transfer of units from the assembly department to the painting department (weighted-average method.)
Note: Enter debits before credits.
Transaction General Journal Debit Credit
1
Answer #1

Equivalent unit of material = 10000+(3300*80%) = 12640

Equivalent unit of conversion = 10000+(3300*30%) = 10990

Cost per equivalent unit of material = (896+13008)/12640 = 1.1

Cost per equivalent unit of conversion = (1800+14685)/10990 = 1.5

Journal entry :

Date account and explanation debit credit
Work in process-Painting department (1.1+1.5)*10000 26000
Work in process-Assembly department 26000
(To record units transferred out)